A company knows by experience that usually shipment , 1/8 of the goods are usually damaged . if it wants to receive 560 pieces of undamaged goods ,how many total goods must it order for shipment ?
Answer by addingup(3677)

8/8-1/8 = 7/8
7/8x = 560
x = 560(8/7) = 640